OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. 12 + 42 + 72 + ... + (3n - 2)2 =

OpenStudy (zzr0ck3r):

=?

OpenStudy (zzr0ck3r):

this makes no sense...(3n-2)^2 would be 1+16...

OpenStudy (anonymous):

this is what it actually looks like sorry 1^2 + 4^2 + 7^2 + ... + (3n - 2)^2 =

OpenStudy (zzr0ck3r):

correct

OpenStudy (zzr0ck3r):

=?

OpenStudy (anonymous):

that is the whole problem...

OpenStudy (zzr0ck3r):

it makes no sense.

OpenStudy (zzr0ck3r):

there is nothing to prove

OpenStudy (zzr0ck3r):

what statement is true?

OpenStudy (zzr0ck3r):

look again...

OpenStudy (anonymous):

how do I prove the problem is true??

OpenStudy (zzr0ck3r):

THERE IS NOTHING TO PROVE you wrote this prove that something = but you didnt tell us what we need to prove its equal to

OpenStudy (zzr0ck3r):

thats like me saying Prove that every animal has

OpenStudy (zzr0ck3r):

has what?!

OpenStudy (zzr0ck3r):

there MUST be something after the = sign

OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false.

OpenStudy (zzr0ck3r):

@jbrimson ?

OpenStudy (zzr0ck3r):

dude look at any other problem that tells you to prove something by induction. do they have something after the equal sign?

OpenStudy (zzr0ck3r):

\(1^2+4^2+......+(3n-2)^2=\frac{1}{2}n(6n^2-3n-1)\) I can show you how to prove this...

OpenStudy (zzr0ck3r):

and since I can prove this, there is nothing else that you could put after the equals sign to make it true. Would you like me to do so?

OpenStudy (anonymous):

12 + 42 + 72 + ... + (3n - 2)2 = n(6n^2-3n-1)/2 sorry I see what I left out

OpenStudy (zzr0ck3r):

lol

OpenStudy (zzr0ck3r):

really dude/?

OpenStudy (zzr0ck3r):

@jbrimson if you want help you should stay in the room of the question....as much as I like doing your work for u....

OpenStudy (zzr0ck3r):

tell me if you have questions